simplify (x-2)*-2(x-2)-15 *= squared
Answer by Alan3354(69443)

simplify (x-2)*-2(x-2)-15 *= squared
----------------
Expand everything and collect terms
=(x^2 - 4x + 4) -2x + 4 -15
=x^2 - 6x - 7
BTW, that's (x-7)*(x+1)